# Lepton Flavor-Violating Higgs Decays Mediated by Ultralight Gauge Boson

### v1

@@ 3.2 Higgs decays with $\boldsymbol{\chi}$ on-shell
where the squared amplitude as a function of $s$ and $t$ is given in Appendix B. The kinematic limits are
$t_{\binom{{\rm max}}{{\rm m_{\ell_{i}}n}}}=\frac{1}{4s}\biggl{[}\left(M_{H}^{2}+m_{\chi}^{2}-m_{\ell_{i}}^{2}-m_{\ell_{j}}^{2}\right)^{2}-\left(\lambda^{1/2}[s,m_{\ell_{i}}^{2},m_{\chi}^{2}]\mp\lambda^{1/2}[s,M_{H}^{2},m_{\ell_{j}}^{2}]\right)^{2}\biggr{]}\,,$ (15)
@@ 3.2.2 Angular Observables
We also examined the decays $H\to\ell_{i}\bar{\ell}_{j}\chi$ as functions of lepton energy ($E_{\ell_{i}}$) and an angular variable $\cos{\theta_{\ell_{i}\ell_{j}}}$. Here, $\theta_{\ell_{i}\ell_{j}}$ represents the angle between the momenta of the two leptons in the rest frame of the $\ell_{i}-\chi$ system, where $\vec{q}_{\ell_{i}}+\vec{q}_{\chi}=\vec{0}$. Consequently, we have $|\vec{p}_{H}|=|\vec{q}_{\ell_{j}}|=\sqrt{E_{H}^{2}-M_{H}^{2}}$ and $|\vec{q}_{\ell_{i}}|=|\vec{q}_{\chi}|=\sqrt{E_{\ell_{i}}^{2}-mi^{2}}$, with
@@
$\frac{d^{2}\Gamma(H\to\ell_{i}\bar{\ell}_{j}\chi)}{dE_{\ell_{i}}d\cos{\theta_{\ell_{i}\ell_{j}}}}=$ $\frac{\lambda^{1/2}[\left(E_{\chi}+E_{\ell_{i}}\right)^{2}\,,M_{H}^{2}\,,m_{\ell_{j}}^{2}]}{(2\pi)^{3}16M_{H}^{2}E_{\chi}}(E_{\chi}+E_{\ell_{i}})|\vec{q}_{\ell_{i}}|~{}\overline{|\mathcal{M}_{H\to\ell_{i}\bar{\ell}_{j}\chi}(\cos{\theta_{\ell_{i}\ell_{j}}},E_{\ell_{i}})|^{2}}\,.$ (18)
where $\lambda$ is the Källén function.
@@ Appendix A Appendix: mixing angles and masses in the tree level model
Appendix B Appendix: Squared Amplitude

### v2

@@ 3.2 Higgs decays with $\boldsymbol{\chi}$ on-shell
where the squared amplitude as a function of $s$ and $t$ is given in Appendix C. The kinematic limits are
$t_{\binom{{\rm max}}{{\rm min}}}=\frac{1}{4s}\biggl{[}\left(M_{H}^{2}+m_{\chi}^{2}-m_{\ell_{i}}^{2}-m_{\ell_{j}}^{2}\right)^{2}-\left(\lambda^{1/2}[s,m_{\ell_{i}}^{2},m_{\chi}^{2}]\mp\lambda^{1/2}[s,M_{H}^{2},m_{\ell_{j}}^{2}]\right)^{2}\biggr{]}\,,$ (15)
@@ 3.2.2 Angular Observables
We also examined the decays $H\to\ell_{i}\bar{\ell}_{j}\chi$ as functions of lepton energy ($E_{\ell_{i}}$) and an angular variable $\cos{\theta_{\ell_{i}\ell_{j}}}$. Here, $\theta_{\ell_{i}\ell_{j}}$ represents the angle between the momenta of the two leptons in the rest frame of the $\ell_{i}-\chi$ system, where $\vec{q}_{\ell_{i}}+\vec{q}_{\chi}=\vec{0}$. Consequently, we have $|\vec{p}_{H}|=|\vec{q}_{\ell_{j}}|=\sqrt{E_{H}^{2}-M_{H}^{2}}$ and $|\vec{q}_{\ell_{i}}|=|\vec{q}_{\chi}|=\sqrt{E_{\ell_{i}}^{2}-m_{\ell_{i}}^{2}}$, with
@@
$\frac{d^{2}\Gamma(H\to\ell_{i}\bar{\ell}_{j}\chi)}{dE_{\ell_{i}}d\cos{\theta_{\ell_{i}\ell_{j}}}}=$ $\frac{(E_{\chi}+E_{\ell_{i}})^{2}|\vec{q}_{\ell_{i}}||\vec{q}_{\ell_{j}}|}{(2\pi)^{3}~{}8M_{H}^{3}E_{\chi}}\overline{|\mathcal{M}_{H\to\ell_{i}\bar{\ell}_{j}\chi}(\cos{\theta_{\ell_{i}\ell_{j}}},E_{\ell_{i}})|^{2}}\,.$ (18)
@@ Appendix A Appendix: mixing angles and masses in the tree level model
Appendix B Appendix: Counterterms for the Amplitude of $H\to\ell_{i}\bar{\ell}_{j}$
The amplitude generated by the triangle diagram for the decay $H\to\ell_{i}\bar{\ell}_{j}$ includes UV-divergent terms, necessitating the introduction of counterterms to absorb these divergences. The one-loop amplitudes and the corresponding form factors are initially expressed in terms of Passarino-Veltman functions (see Eq. (3.1)). For numerical calculations, however, we work with the renormalized amplitude, where the divergences are canceled through counterterms. The associated counterterm Lagrangian is given by
$\mathcal{L}_{{\rm CT}}=C_{f_{ij}}\bar{\ell}_{i}\ell_{j}H+C_{g_{ij}}\bar{\ell}_{i}\gamma_{5}\ell_{j}H+{\rm h.c.}\,,$ (25)
where the coefficients of the scalar and pseudoscalar operators are specified as:
$C_{f_{ij}}$ $=\frac{m_{\ell_{k}}}{2m_{\chi}^{2}v\bar{\epsilon}_{{\rm uv}}}\left(M_{H}^{2}-6m_{\ell_{k}}^{2}+6m_{\chi}^{2}\right)(f_{ik}f_{jk}-g_{ik}g_{jk})\,,$
$C_{g_{ij}}$ $=\frac{m_{\ell_{k}}}{2m_{\chi}^{2}v\bar{\epsilon}_{{\rm uv}}}\left(M_{H}^{2}-6m_{\ell_{k}}^{2}+6m_{\chi}^{2}\right)(f_{ik}g_{jk}-f_{jk}g_{ik})\,,$ (26)
with the effective couplings $f_{ij}$ and $g_{ij}$ defined in Eq. (10), and $\frac{1}{\bar{\epsilon}_{{\rm uv}}}\equiv\frac{1}{\epsilon_{{\rm uv}}}-\gamma_{E}+\ln 4\pi$. The amplitude is renormalized using the $\overline{\rm MS}$-scheme, ensuring that only finite contributions remain in the final calculation.
It is important to observe that the counterterm Lagrangian in Eq. (25) comprises two distinct contributions: the first term corresponds to a scalar operator, while the second term represents a pseudoscalar operator. The pseudoscalar term does not contribute to the amplitude in the flavor-conserving scenario. This indicates that quantum corrections primarily generate the scalar operator, similar to the behavior observed in the Standard Model.
Appendix C Appendix: Squared Amplitude